OpenStudy (anonymous):

what is this standard equation of an ellipse, in y-intercept (y=???) form? (x+4)^2 \6 + (y-9)^2 \2 = 1

OpenStudy (amistre64):

multiply by 2 and then by 6 2(x+4)^2 + 6(y-9)^2 = 12 ; subtract that left part 6(y-9)^2 = 12 - 2(x+4)^2 ; divide by 6 (y-9)^2 = 12 -2(x+4)^2//6 ; square both sides y-9 = sqrt(12 -2(x+4)^2) ; add 9 y = sqrt(12 -2(x+4)^2) +9

OpenStudy (amistre64):

\[y = \sqrt{12-2(x+4)^2}+9\]

OpenStudy (amistre64):

thats the best I can do with it...maybe add a +- to it, but not much else.

OpenStudy (anonymous):

thank you !!

OpenStudy (amistre64):

denada